A 800 kg drag racing car gets up to 100 miles per hour from rest in 0.8 s.

a. What is the change in momentum of the car?

b. What is the average force exerted on the car?

c. What exerts that force?
Answer #1

here,

the mass of car , m = 800 kg

the initial speed of car , u = 100 mph

u = 44.7 m/s

time taken , t = 0.8 s

a)

the change in momentum of the car , dP = Pf - Pi

dP = m * 0 - m * u

dP = - 800 * 44.7 kg.m/s

dP = - 3.58 * 10^4 kg.m/s

b)

the average force exerted on the car , F = dP /t

F = - 3.58 * 10^4 /(0.8) N

F = - 4.47 * 10^4 N

c)

this force is exerted by road on the car